# Quantitative description of strongly correlated materials by combining downfolding techniques and tensor networks

### v1

@@ III.1 Trans-polyacetylene
III.1.1 Downfolding
The first step in this approach is the DFT simulation. Fig. 3 presents the band structure for the three tPA geometries. The valence and conduction bands, predominantly of carbon character, are separated by a small band gap. As expected, the gap increases with a larger BLA, however, even for tPA3, its value remains significantly underestimated. From these bands, two MLWFs are constructed, each centered on a carbon atom within the unit cell. Consequently, the effective model corresponds to a two-band Hubbard model at half-filling.
The hopping parameters , obtained via Hartree correction of the tight-binding parameters, and the direct Coulomb interaction from cRPA are presented in Table 1 for tPA3. The remaining parameters, as well as those for the other two structures, are provided in Appendix C. By subtracting a constant chemical potential, the diagonal terms of the on-site hopping matrix are adjusted to set the smallest value on the diagonal to zero.
@@ III.1.1 Downfolding
For each geometry, we analyze multiple models. The hopping range is selected to include all terms listed in Table 1 except the parameter above the diagonal of the matrix, , ensuring consistency with the corresponding physical distances. Regarding interactions, we incorporate all terms up to a specified range . For , only interactions between neighboring orbitals are included; for , next-to-nearest neighbor interactions are added, and so on. We extend this up to , which includes all direct interactions larger than 2 eV in Table 1, for example. As previously noted, all interaction terms within this range are considered, encompassing not only the direct interactions but also the exchange terms , those of the form , and even interactions involving three or four different orbitals, such as and . Parameters smaller than 10 meV are excluded.
III.1.2 Model solution
The models defined above are implemented and solved using tensor network code, as described in Section V. With the quasiparticle Ansatz, the first single-particle and single-hole excitation energies, and respectively, are calculated for momenta along the chain. The minimum of the sum of these excitations determines the band gap of the system
@@ III.1.2 Model solution
Fig. 4 illustrates the band gap for different geometries as a function of the range . The dashed line represents the experimental value reported by Wang et al. [91], obtained via scanning tunneling spectroscopy and differential conductance spectroscopy on single tPA chains. The dotted line marks the value computed using equation-of-motion CCSD (EOM-CCSD), extrapolated to the thermodynamic limit by Moerman et al. [123].
@@
Table 2 compares the band gaps predicted by our method (DFT+TN) with those obtained using other computational methods performed with VASP. Regardless of the chosen functional, DFT tends to underestimate the band gap. Similarly, partially self-consistent calculations generally yield smaller values. Our results align well with fully self-consistent calculations but tend to be slightly larger.
@@
III.2 Polythiophene
III.2.1 Downfolding
PT, illustrated in Fig. 1b, has a more complex structure than tPA. Again, we start from a DFT calculation to obtain the Kohn-Sham eigenvalues. From the resulting band structure (see Fig. 5), a band gap of 1.1 eV is observed at the -point. The orbitals exhibit noticeable hybridization, leading to electron delocalization. To construct sufficiently localized Wannier functions, we focus on the target space spanned by the six bands associated with two sulfur and four carbon atoms. As shown in Fig. 5, the Wannier-interpolated bands accurately reproduce the four highest valence bands and the two lowest conduction bands. The effective model consists of six bands, a relatively large number compared to typical tight-binding or low-energy models. This results in a filling of eight electrons distributed among six orbitals, highlighting the complexity of the system.
@@ III.2.2 Model solution
From the six-band Hubbard model described above, a band gap of 2.94 eV is obtained at momentum , with convergence achieved at bond lengths above 300 (see Appendix D). This result shows good agreement with the single-shot calculation (see Table 4), while fully self-consistent predicts a value 0.65 eV higher. Although the effective model neglects exchange and longer-range interactions, our methodology aligns more closely with calculations than hybrid DFT functionals such as HSE06 [125] and B3LYP. This suggests that even with its simplifications, the model captures essential electronic properties more effectively than commonly used exchange-correlation functionals. That said, the delocalized nature of the electrons and relatively weak electron correlation make this material less ideal for this approach, as strong localization effects typically enhance the accuracy of such models. Nonetheless, we demonstrate that it remains effective in describing the system’s fundamental electronic features.
III.3 Sr2CuO3
III.3.1 Downfolding
The quasi-1D material Sr2CuO3, shown in Fig. 1c, contains 1D CuO3 chains. DFT calculations in VASP on a -mesh produce the band structure given in Fig. 6. The bands crossing the Fermi energy originate from copper orbitals. A supercell containing four copper atoms was necessary to access nearest-neighbor interaction parameters in VASP, so that four such copper bands are present in the band structure. Single chains can then be considered as one-band Hubbard models at half-filling. The derived model parameters for Sr2CuO3 can be found in Table 5 and are all taken into account. The on-site hopping term is a constant chemical potential, which does not affect the physical properties since the number of electrons is fixed. Due to technical complications (see Section VB), the Hartree double-counting corrections (13) cannot be determined and are therefore omitted in this case. The direct interactions only contribute to a shift in the chemical potential, which is not physically relevant. Since both and are already small, the correction terms are expected to be negligible, justifying the approximation.
@@ III.3.1 Downfolding
III.3.2 Model solution
In Fig. 7, the staggered magnetization, , of the ground state is plotted as a function of the bond dimension. At first glance, it appears to converge toward a small but nonzero value, suggesting the presence of antiferromagnetism within the chain. However, the existence of antiferromagnetic order in a purely 1D -invariant spin chain is ruled out by the Hohenberg-Mermin-Wagner theorem [130, 131, 132], or more precisely, by its zero-temperature extension [133]. It is well known, however, that MPS simulations tend to develop such order and break the symmetry as a means of reducing entanglement, though the symmetry should be restored in the limit of increasing bond dimension. The experimentally observed antiferromagnetism must therefore be a consequence of spin-orbit or interchain coupling, and the precise value of the order parameter cannot be predicted with the current approach.
@@ IV Conclusion and outlook
V Computational details
V.1 Structure relaxation
Geometry relaxation is performed using VASP [137, 138] with the projector augmented-wave method [139]. For tPA1 and tPA2, the PBE [99] and B3LYP [21] functionals are used, respectively, with a 20 Å vacuum layer to isolate the chains. A plane-wave energy cutoff of 600 eV is applied on a -grid along the chain axis. Electronic minimization proceeds until the energy difference falls below eV, while structural relaxation continues until forces converge to eV. See Appendix A for the resulting energy surface.
V.2 Downfolding
Downfolding for tPA and PT is performed via 1D DFT calculations in PySCF [140], employing the PBE functional. For tPA, the all-electron correlation-consistent polarized valence triple-zeta (cc-pVTZ) basis set [141] is used, while for PT, the pseudopotential GTH-DZVP basis from the CP2K software package [142] is applied. MLWFs are constructed using Wannier90 [143], and interaction and hopping parameters are extracted via an in-house cRPA implementation [136].
For tPA, the initial DFT calculation is performed with 28 -points along the chain to ensure parameter convergence, while for PT, 12 -points are used. The direct interactions exhibit a slow convergence with respect to the number of -points, following a dependence (see Appendix B). This behavior arises from the Coulomb divergence in the Gaussian density fitting scheme implemented in PySCF [144, 145]. As proposed by Berkelbach and coworkers [145], we extrapolated these parameters based on a reliable fit. The hopping and other interaction parameters do not suffer from this issue.
V.3 Tensor network calculations
Once the Hamiltonian is constructed, infinite two-site DMRG is performed to compute the ground state of the system while dynamically determining the bond dimension, followed by several VUMPS iterations to guarantee convergence. The relative tolerance is set to . Excitations on top of the ground state are determined using the quasiparticle algorithm, as introduced by Haegeman et al. [115]. We look for the first particle and hole excitation as a function of the momentum and identify their sum as the required energy to create a free electron-hole pair. The band gap of the model is then extracted as the minimum. Excitons, on the other hand, correspond to excitations in the trivial sector.
@@ Acknowledgements.
- Hybertsen et al. [1989] M. S. Hybertsen, M. Schlüter, and N. E. Christensen, Calculation of coulomb-interaction parameters for using a constrained-density-functional approach, Phys. Rev. B 39, 9028 (1989).
@@
- Anisimov et al. [1997] V. I. Anisimov, F. tiawan, and A. I. Lichtenstein, First-principles calculations of the electronic structure and spectra of strongly correlated systems: the LDA+U method, Journal of Physics: Condensed Matter 9, 767 (1997).
@@
- Hedin and Lundqvist [1970] L. Hedin and S. Lundqvist, Effects of electron-electron and electron-phonon interactions on the one-electron states of solids (Academic Press, 1970) pp. 1–181.
- [25] F. Aryasetiawan and F. Nilsson, Downfolding methods in many-electron theory.
@@
- Verstraete and Cirac [2004a] F. Verstraete and J. I. Cirac, Renormalization algorithms for quantum-many body systems in two and higher dimensions (2004a), arXiv:cond-mat/0407066 [cond-mat.str-el] .
@@
- Verstraete and Cirac [2004b] F. Verstraete and J. I. Cirac, Valence-bond states for quantum computation, Phys. Rev. A 70, 060302 (2004b).
@@
- Kaltak [2015] M. Kaltak, Merging GW with DMFT, Ph.D. thesis, Vienna (2015).
@@
- Perez-Garcia et al. [2007] D. Perez-Garcia, F. Verstraete, M. M. Wolf, and J. I. Cirac, Matrix product state representations (2007), arXiv:quant-ph/0608197 [quant-ph] .
@@
- Windom et al. [2024] Z. W. Windom, A. Lam, A. Perera, and R. J. Bartlett, An assessment of frozen natural orbitals and band gaps using equation of motion coupled cluster theory: a case study on polyacene and trans-polyacetylene (2024), arXiv:2402.08776 [physics.chem-ph] .
@@ B.2 Polythiophene
The convergence as a function of -points for some relevant parameters of the Hubbard model of polythiophene (PT) is shown in Fig. 10. For the direct interactions, fits are made to extrapolate to infinite . These fits exclude the first three data points (). Some of the exchange and parameters are not entirely converged at , but these are not considered in the effective model.
@@ Appendix C Parameters trans-polyacetylene
We report the parameters included in the effective models for the three geometries of tPA.

### v2

@@ IV Results and Discussion
IV.1.1 Downfolding
The first step in this approach is the DFT simulation. Figure 3 presents the band structure for the three tPA geometries. The valence and conduction bands, predominantly of carbon character, are separated by a small band gap. As expected, the gap increases with a larger BLA, however, even for tPA3, its value remains significantly underestimated. From these bands, two MLWFs are constructed, each centered on a carbon atom within the unit cell. Consequently, the effective model corresponds to a two-band Hubbard model at half-filling.
The hopping parameters , obtained via Hartree correction of the tight-binding parameters, and the direct Coulomb interaction from cRPA are presented in Table 1 for tPA3. The remaining parameters, as well as those for the other two structures, are provided in the Supporting Information. By subtracting a constant chemical potential, the diagonal terms of the on-site hopping matrix are adjusted to set the smallest value on the diagonal to zero.
@@ IV.1.1 Downfolding
For each geometry, we analyze multiple models. The hopping range is selected to include all terms listed in Table 1 except the parameter above the diagonal of the matrix, , ensuring consistency with the corresponding physical distances. Regarding interactions, we incorporate all terms up to a specified range . For , only interactions between neighboring orbitals are included; for , next-nearest-neighbor interactions are added, and so on (see Figure 4a). We extend this up to , which includes all direct interactions larger than 2 eV in Table 1, for example. As previously noted, all interaction terms within this range are considered, including the direct , the exchange , and the bond-charge interactions (or correlated hopping) [145, 146], which have been previously studied in the context of polyacetylene [147]. We find that the pair-hopping parameters are identical to the exchange terms, i.e. , and are always included whenever exchange is considered. Additionally, interactions involving three or four different orbitals, such as and , are also taken into account. Interaction parameters smaller than 10 meV are excluded.
IV.1.2 Model Solution
The models defined above are implemented and solved using tensor network code, as described in Section III. With the quasiparticle Ansatz, the first single-particle and single-hole excitation energies, and respectively, are calculated for momenta along the chain. The minimum of the sum of these excitations determines the band gap of the system
@@ IV.1.2 Model Solution
Figure 4b illustrates the band gap for different geometries as a function of the range . The dashed line represents the experimental value reported by Wang et al. [98], obtained via scanning tunneling spectroscopy and differential conductance spectroscopy on single tPA chains. The dotted line marks the value computed using equation-of-motion CCSD (EOM-CCSD), extrapolated to the thermodynamic limit by Moerman et al. [143].
@@
Table 2 compares the band gaps predicted by our method (DFT+TN) with those obtained using other computational methods performed with VASP. The vacuum between chains was reduced to 12.5 Å in the calculations. Regardless of the chosen functional, DFT tends to underestimate the band gap. Similarly, partially self-consistent calculations generally yield smaller values. Our results align well with fully self-consistent calculations but tend to be slightly larger.
@@
IV.2 Polythiophene
IV.2.1 Downfolding
PT, illustrated in Figure 1b, has a more complex structure than tPA. Again, we start from a DFT calculation to obtain the Kohn-Sham eigenvalues. From the resulting band structure (see Figure 5), a band gap of 1.1 eV is observed at the -point. The orbitals exhibit noticeable hybridization, leading to electron delocalization. To construct sufficiently localized Wannier functions, we focus on the target space spanned by the six bands associated with two sulfur and four carbon atoms. As shown in Figure 5, the Wannier-interpolated bands accurately reproduce the four highest valence bands and the two lowest conduction bands. The effective model consists of six bands, a relatively large number compared to typical tight-binding or low-energy models. This results in a filling of eight electrons distributed among six orbitals, highlighting the complexity of the system.
@@ IV.2.2 Model solution
From the six-band Hubbard model described above, a band gap of 2.94 eV is obtained at momentum , with convergence achieved at bond lengths above 300. This result shows good agreement with the single-shot calculation (see Table 4), while fully self-consistent predicts a value 0.65 eV higher. Although the effective model neglects exchange and longer-range interactions, our methodology aligns more closely with calculations than hybrid DFT functionals such as HSE06 [148] and B3LYP. This suggests that even with its simplifications, the model captures essential electronic properties more effectively than commonly used exchange-correlation functionals. That said, the delocalized nature of the electrons and relatively weak electron correlation make this material less ideal for this approach, as strong localization effects typically enhance the accuracy of such models. Nonetheless, we demonstrate that it remains effective in describing the system’s fundamental electronic features.
IV.3 Sr2CuO3
IV.3.1 Downfolding
The quasi-1D material Sr2CuO3, shown in Figure 1c, contains 1D CuO3 chains. DFT calculations in VASP on a -mesh produce the band structure given in Figure 6. The bands crossing the Fermi energy originate from copper orbitals. A supercell containing four copper atoms was necessary to access nearest-neighbor interaction parameters in VASP, so that four such copper bands are present in the band structure. Single chains can then be considered as one-band Hubbard models at half-filling. The derived model parameters for Sr2CuO3 can be found in Table 5 and are all taken into account. The on-site hopping term is a constant chemical potential, which does not affect the physical properties since the number of electrons is fixed. Due to technical complications (see Section III.2), the Hartree double-counting corrections (13) cannot be determined and are therefore omitted in this case. The direct interactions only contribute to a shift in the chemical potential, which is not physically relevant. Since both and are already small, the correction terms are expected to be negligible, justifying the approximation.
@@ IV.3.1 Downfolding
IV.3.2 Model Solution
In Figure 7, the staggered magnetization, , of the ground state is plotted as a function of the bond dimension. At first glance, it appears to converge toward a small but nonzero value, suggesting the presence of antiferromagnetism within the chain. However, the existence of antiferromagnetic order in a purely 1D -invariant spin chain is ruled out by the Hohenberg-Mermin-Wagner theorem [153, 154, 155], or more precisely, by its zero-temperature extension [156]. It is well known, however, that MPS simulations tend to develop such order and break the symmetry as a means of reducing entanglement, though the symmetry should be restored in the limit of increasing bond dimension. The experimentally observed antiferromagnetism must therefore be a consequence of spin-orbit or interchain coupling, and the precise value of the order parameter cannot be predicted with the current approach.
@@ Acknowledgements.
- Anisimov et al. [1997] V. I. Anisimov, F. Aryasetiawan, and A. I. Lichtenstein, First-principles calculations of the electronic structure and spectra of strongly correlated systems: the LDA+U method, Journal of Physics: Condensed Matter 9, 767 (1997).
@@
- Hedin and Lundqvist [1970] L. Hedin and S. Lundqvist, Effects of Electron-Electron and Electron-Phonon Interactions on the One-Electron States of Solids, edited by F. Seitz, D. Turnbull, and H. Ehrenreich, Solid State Physics, Vol. 23 (Academic Press, 1970) pp. 1–181.
- Aryasetiawan and Nilsson [2022] F. Aryasetiawan and F. Nilsson, Downfolding Methods in Many-Electron Theory (AIP Publishing LLC, 2022).
@@
- Pham et al. [2020] H. Q. Pham, M. R. Hermes, and L. Gagliardi, Periodic electronic structure calculations with the density matrix embedding theory, Journal of Chemical Theory and Computation 16, 130 (2020).
@@
- Nusspickel and Booth [2022] M. Nusspickel and G. H. Booth, Systematic improvability in quantum embedding for real materials, Phys. Rev. X 12, 011046 (2022).
@@
- Baiardi and Reiher [2020] A. Baiardi and M. Reiher, The density matrix renormalization group in chemistry and molecular physics: Recent developments and new challenges, The Journal of Chemical Physics 152, 040903 (2020).
- Barcza et al. [2013] G. Barcza, W. Barford, F. Gebhard, and Ö. Legeza, Excited states in polydiacetylene chains: A density matrix renormalization group study, Phys. Rev. B 87, 245116 (2013).
- Fertitta et al. [2014] E. Fertitta, B. Paulus, G. Barcza, and Ö. Legeza, Investigation of metal–insulator-like transition through the ab initio density matrix renormalization group approach, Phys. Rev. B 90, 245129 (2014).
- Timár et al. [2016] M. Timár, G. Barcza, F. Gebhard, L. Veis, and Ö. Legeza, Hückel–hubbard–ohno modeling of -bonds in ethene and ethyne with application to trans-polyacetylene, Physical Chemistry Chemical Physics 18, 18835 (2016).
- [81] M. A. Werner, A. Menczer, and Ö. Legeza, Tensor network state methods and quantum information theory for strongly correlated molecular systems, arXiv (Condensed Matter.Strongly Correlated Electrons), 2025-01-30, 10.48550/arXiv.2402.08776 (accessed 2025-02-18).
- Alvertis et al. [2025] A. M. Alvertis, A. Khan, and N. M. Tubman, Compressing hamiltonians with ab initio downfolding for simulating strongly-correlated materials on quantum computers, Phys. Rev. Appl. 23, 044028 (2025).
- [83] F. Verstraete and J. I. Cirac, Renormalization algorithms for quantum-many body systems in two and higher dimensions, arXiv (Condensed Matter.Strongly Correlated Electrons), 2004-07-02, 10.48550/arXiv.cond-mat/0407066 (accessed 2025-02-17).
@@
- Verstraete and Cirac [2004] F. Verstraete and J. I. Cirac, Valence-bond states for quantum computation, Phys. Rev. A 70, 060302 (2004).
@@
- Kaltak [2015] M. Kaltak, Merging GW with DMFT, Ph.D. thesis, Universität Wien, Vienna (2015).
@@
- Zhu et al. [2020] T. Zhu, Z.-H. Cui, and G. K.-L. Chan, Efficient formulation of ab initio quantum embedding in periodic systems: Dynamical mean-field theory, Journal of Chemical Theory and Computation 16, 141 (2020).
@@
- [117] D. Perez-Garcia, F. Verstraete, M. M. Wolf, and J. I. Cirac, Matrix product state representations, arXiv (Quantum Physics), 2006-08-25, 10.48550/arXiv.quant-ph/0608197 (accessed 2025-02-11).
@@
- [142] Z. W. Windom, A. Lam, A. Perera, and R. J. Bartlett, An assessment of frozen natural orbitals and band gaps using equation of motion coupled cluster theory: a case study on polyacene and trans-polyacetylene, arXiv (Physics.Chemical Physics), 2024-02-13, 10.48550/arXiv.2402.08776 (accessed 2025-02-18).
- Schubin et al. [1934] S. Schubin, S. Wonsowsky, and R. H. Fowler, On the electron theory of metals, Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character 145, 159 (1934).
- Hirsch [1989] J. Hirsch, Bond-charge repulsion and hole superconductivity, Physica C: Superconductivity and its Applications 158, 326 (1989).
- Kivelson et al. [1987] S. Kivelson, W.-P. Su, J. R. Schrieffer, and A. J. Heeger, Missing bond-charge repulsion in the extended hubbard model: Effects in polyacetylene, Phys. Rev. Lett. 58, 1899 (1987).
@@ B.2 Polythiophene
The convergence as a function of -points for some relevant parameters of the Hubbard model of polythiophene (PT) is shown in Fig. 12. For the direct interactions, fits are made to extrapolate to infinite . These fits exclude the first three data points (). Some of the exchange and bond-charge parameters are not entirely converged at , but these are not considered in the effective model.
@@ Appendix C Parameters trans-Polyacetylene
We report the parameters included in the effective models for the three geometries of tPA. The interaction parameters can also be found in csv format in the GitHub repository [126].